OpenStudy (anonymous):

Describe the pattern in the numbers and write the next two numbers 0.0625,o.125,0.25

OpenStudy (mr.math):

Hint: \[0.0625=\frac{1}{16}\] \[0.125=\frac{1}{8}\] \[0.25=\frac{1}{4}\]

OpenStudy (anonymous):

a number is formed by multiply the previous one by 2

OpenStudy (anonymous):

the next two numbers in the series are 0.50 and 1 or as a fraction 1/2 and 1